OpenStudy (anonymous):

how do i factor x^2-9x+10?

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to 10 and add to -9

OpenStudy (anonymous):

so -10 and 1?

jimthompson5910 (jim_thompson5910):

notice how -10 times 1 = -10 though

jimthompson5910 (jim_thompson5910):

but we want them to multiply to +10 or just 10

OpenStudy (anonymous):

-10 and -1? but then doesnt that equal -11?

jimthompson5910 (jim_thompson5910):

yes so we get +10 when they multiply but -11 when they add

OpenStudy (anonymous):

so then how do you get -9?

OpenStudy (anonymous):

the only other factors besides 1 and 10 are 5 and 2 but that equals 7

jimthompson5910 (jim_thompson5910):

well it turns out that there are no two whole numbers that both a) multiply to 10 AND b) add to -9 at the same time so this all means that x^2-9x+10 CANNOT be factored

jimthompson5910 (jim_thompson5910):

so we say x^2-9x+10 is prime

OpenStudy (anonymous):

oh! ok so next i have 4y^2-49 how would i do that?

jimthompson5910 (jim_thompson5910):

hint: difference of squares

OpenStudy (anonymous):

so would i need to find the square root of both numbers or just one?

OpenStudy (anonymous):

sorry im horrible at math... :/

jimthompson5910 (jim_thompson5910):

of both, so what is the square root of 4y^2?

OpenStudy (anonymous):

the square root of 8 is 2.82

OpenStudy (anonymous):

and 49 is 7

jimthompson5910 (jim_thompson5910):

4y^2 is the same as (2y)^2 you take the square root of 4 to get 2 you take the square root of y^2 to get y so that's how I got 2y

jimthompson5910 (jim_thompson5910):

so 4y^2 - 49 turns into (2y)^2 - 7^2

jimthompson5910 (jim_thompson5910):

now you use the difference of squares formula

OpenStudy (anonymous):

so it would be 2^2-7^2?

jimthompson5910 (jim_thompson5910):

the difference of squares formula is a^2 - b^2 = (a-b)(a+b)

jimthompson5910 (jim_thompson5910):

so in this case a = 2y b = 7

OpenStudy (anonymous):

oh! so it's (2-7) (2+7)

jimthompson5910 (jim_thompson5910):

very close, but you're missing the y

jimthompson5910 (jim_thompson5910):

it should be (2y-7) (2y+7)

OpenStudy (anonymous):

ok! im sorry but i have two more i dont understand. can you help me real quick?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

ok so there's 4n^2-20n+25

jimthompson5910 (jim_thompson5910):

have you learned about perfect square trinomials?

OpenStudy (anonymous):

my teacher just went over what a trinomial was but i dont know what to do.

OpenStudy (anonymous):

would you just find the square root of 4 and 25?

jimthompson5910 (jim_thompson5910):

yes what is the square root of each?

OpenStudy (anonymous):

2 and 5

jimthompson5910 (jim_thompson5910):

notice how 2*5 = 10 and if you double that, you get 2*10 = 20

jimthompson5910 (jim_thompson5910):

because this is equal to the middle coefficient (ignore the negative), this means we have a perfect square trinomial

OpenStudy (anonymous):

so it would be (2+5) (2+5)?

OpenStudy (anonymous):

i mean (2n+5)(2n+5)

jimthompson5910 (jim_thompson5910):

close, but the middle coefficient is negative, so you have to change all the + signs to - signs so it should be (2n-5)(2n-5)

OpenStudy (anonymous):

ok and now there's 100-w^2

OpenStudy (anonymous):

would this be common factoring? because i have to also place them in either trinomial, difference of squares or common factoring

jimthompson5910 (jim_thompson5910):

this is another difference of squares

jimthompson5910 (jim_thompson5910):

100 is a perfect square (100 = 10^2) w^2 is a perfect square

OpenStudy (anonymous):

so it would be (10-w)(10-w)?

jimthompson5910 (jim_thompson5910):

yep, you nailed it

jimthompson5910 (jim_thompson5910):

oh wait, sorry

jimthompson5910 (jim_thompson5910):

it's actually (10-w)(10+w) the signs should alternate

OpenStudy (anonymous):

wow i suck haha

jimthompson5910 (jim_thompson5910):

since it's a difference of squares and the rule is a^2 - b^2 = (a-b)(a+b)

jimthompson5910 (jim_thompson5910):

no you don't, get that idea out of your head, it's definitely not true and not helpful

jimthompson5910 (jim_thompson5910):

you just made a silly typo

OpenStudy (anonymous):

so it's (w-10)(w+10)?

jimthompson5910 (jim_thompson5910):

no you just needed to change one sign so it's (10-w)(10+w)

jimthompson5910 (jim_thompson5910):

100 - w^2 = 10^2 - w^2 100 - w^2 = (10-w)(10+w)

OpenStudy (anonymous):

ok so how do i figure out if one of them is common factoring because i also have 2x^2-8x, x^2-9x, 2m^2-9m+4, and 4x^2+x

OpenStudy (anonymous):

if you don't know, that's fine

jimthompson5910 (jim_thompson5910):

2x^2 and -8x have a 2x in common, so you factor that out to get 2x^2 - 8x = 2x(x - 4)

OpenStudy (anonymous):

so 2x(x-4) would be the answer?

jimthompson5910 (jim_thompson5910):

yes after factoring 2x^2-8x

jimthompson5910 (jim_thompson5910):

notice how if you distributed that 2x back through, you'd get 2x^2-8x back again

OpenStudy (anonymous):

oh so 2x(x-4) is the factored answer?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

im sorry i keep asking you questions but for 2m^2-9m+4 would i just find the common factor of 2 and 4?

jimthompson5910 (jim_thompson5910):

hmm this is pretty complicated, so one sec

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

ok the steps are pretty long and drawn out, so I had to post them elsewhere (I was able to post a table and use different colors and such, so that's why I posted it there) anyways, it's posted here http://www.algebra.com/algebra/homework/equations/Equations.faq.question.787996.html let me know what you think and feel free to ask questions on any step you have

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

you're a genius!

jimthompson5910 (jim_thompson5910):

thanks lol

jimthompson5910 (jim_thompson5910):

sounds like things are clicking now, which is great

OpenStudy (anonymous):

you should literally be my algebra 2 teacher. i learn more from you hahaha

jimthompson5910 (jim_thompson5910):

I'm glad you are learning, but I'm sure you learned stuff from your teacher too

OpenStudy (anonymous):

im pretty sure i can do the last two by myself but i just have one last quick question. would x^2-9x and 4x^2+x be difference of squares or common factoring?

jimthompson5910 (jim_thompson5910):

x^2-9x would be common factoring (they have x in common) so is 4x^2 + x (this also has x in common)

OpenStudy (anonymous):

so x^2-9x would be x(x-3)?

jimthompson5910 (jim_thompson5910):

close, but it's really x^2 - 9x = x(x-9)

jimthompson5910 (jim_thompson5910):

to check, distribute that x back through and you will get x^2 - 9x again

OpenStudy (anonymous):

oh ok i get it now. so that makes 4x^2+x= 4x(x+x)?

jimthompson5910 (jim_thompson5910):

4x^2 and x do NOT have 4 in common

OpenStudy (anonymous):

they have x in common so it would be x(x+4x)?

jimthompson5910 (jim_thompson5910):

actually it should be 4x^2 + x = x(4x + 1)

OpenStudy (anonymous):

is that because having the x where the 1 is would make it x^2?

OpenStudy (anonymous):

instead of just x?

jimthompson5910 (jim_thompson5910):

well it's because x is also x*1

jimthompson5910 (jim_thompson5910):

so that means 4x^2 + x x*4x + x*1 x(4x + 1)

jimthompson5910 (jim_thompson5910):

see how that works?

OpenStudy (anonymous):

yup! well that's all of my homework for tonight. thank you so much again! you don't understand how much you've helped me :)

jimthompson5910 (jim_thompson5910):

you're welcome, I'm glad to be of help